# Reasoning to Edit: Hypothetical Instruction-Based Image Editing with Visual Reasoning

### v1

@@ 4 Experiments
4.1 Setups: Datasets, Metrics, and Details
Datasets: We use Reason50K for both training and evaluation. Specifically, for each reasoning category, samples are randomly selected for validation, while the remaining samples are used for training. In addition, we assess the reasoning capability of ReasonBrain on two external benchmarks: ReasonEdit [17] and EditWorld [46]. To evaluate generalization on conventional understanding scenarios, we further test on the MagicBrush Test Set [48] and the Emu Edit Test Set [38]. Metrics: To evaluate performance under reasoning scenarios, we adopt three metrics: CLIP Score [32], MLLM Score [46], and Instruction Alignment (Ins-Align) [17]. Specifically, we engage ten annotators to manually evaluate the outputs produced by different methods on the ReasonEdit++ dataset. The primary criterion is how well the edited image aligns with the given instruction. We then average the scores from all annotators to compute the final result. For understanding scenarios, we adopt L1 distance, CLIP image similarity, DINO similarity, CLIP text-image similarity, and CLIP text-image direction similarity as evaluation metrics. Implementation Details: During training, we adopt the pre-trained LLaVAv1.1-7B [27] and QFormer [22] and employ DeepSpeed [3] Zero-2 to perform LoRA [14] fine-tuning, with rank and alpha of 8 and 16, respectively. Following [17, 7], we expand the original LLM vocabulary with 32 new tokens, and the QFormer is composed of 6 transformer layers and 77 learnable query tokens. For the base editing model, we implement it with Flux [21] using FLUX.1-dev, which consists of 12B parameters. Models for other qualitative results are implemented using SD series [6, 39, 35, 2] with their original codebase. Our model is trained with a batch size of 16, and we use AdamW [28] optimizer to train the model with the weight decay as 1e-2 and the learning rate as 1e-3. All the experiments are conducted on 16 H20 GPUs. For a fair comparison, all baseline models are fine-tuned on the same training set used by ReasonBrain.
@@ 4.2 Quantitative Results
| InstructPix2Pix [5] | 0.083 | 0.825 | 0.211 | 0.207 | 0.846 | 0.678 | 0.153 | 0.785 | 0.191 | 0.196 | 0.628 | 0.225 | 0.160 | 0.771 | 0.326 |
| MagicBrush [48] | 0.102 | 0.844 | 0.335 | 0.228 | 0.877 | 0.725 | 0.162 | 0.802 | 0.344 | 0.209 | 0.635 | 0.359 | 0.175 | 0.790 | 0.441 |
| MGIE [7] | 0.098 | 0.802 | 0.288 | 0.213 | 0.832 | 0.685 | 0.155 | 0.761 | 0.328 | 0.201 | 0.622 | 0.322 | 0.167 | 0.754 | 0.406 |
| SmartEdit [17] | 0.118 | 0.849 | 0.602 | 0.226 | 0.881 | 0.779 | 0.165 | 0.823 | 0.385 | 0.211 | 0.655 | 0.361 | 0.180 | 0.802 | 0.532 |
| UltraEdit [51] | 0.156 | 0.861 | 0.584 | 0.231 | 0.922 | 0.826 | 0.193 | 0.869 | 0.482 | 0.209 | 0.669 | 0.458 | 0.197 | 0.830 | 0.588 |
| PixWizard [25] | 0.161 | 0.881 | 0.481 | 0.234 | 0.951 | 0.833 | 0.132 | 0.863 | 0.393 | 0.172 | 0.697 | 0.389 | 0.175 | 0.848 | 0.524 |
@@
| InstructPix2Pix [5] | 0.078 | 0.834 | 0.219 | 0.121 | 0.762 | 0.115 | 0.837 | 0.245 | 0.093 | 0.767 |
| MagicBrush [48] | 0.090 | 0.838 | 0.222 | 0.100 | 0.776 | 0.123 | 0.883 | 0.261 | 0.058 | 0.871 |
| MGIE [7] | 0.083 | 0.746 | 0.231 | 0.163 | 0.594 | 0.116 | 0.745 | 0.251 | 0.162 | 0.577 |
| SmartEdit [17] | 0.092 | 0.858 | 0.274 | 0.119 | 0.771 | 0.119 | 0.895 | 0.262 | 0.094 | 0.820 |
| Emu Edit [38] | 0.109 | 0.859 | 0.231 | 0.094 | 0.819 | 0.135 | 0.897 | 0.261 | 0.052 | 0.879 |
| UltraEdit [51] | 0.107 | 0.844 | 0.283 | 0.071 | 0.793 | - | 0.868 | - | 0.088 | 0.792 |
| PixWizard [25] | 0.104 | 0.845 | 0.248 | 0.069 | 0.798 | 0.124 | 0.884 | 0.265 | 0.063 | 0.876 |
@@ 4.3 Qualitative Results
As shown in Fig. 6, we visualize editing results of ReasonBrain and selected SOTA methods across four different reasoning scenarios. Our method demonstrates superior capability in executing hypothetical editing instructions by accurately reasoning about user intent, affected objects, and their plausible state transitions, while also maintaining stability in non-edited regions. For instance, in the first row (physical reasoning), our method successfully generates a physically plausible and contextually coherent scene–depicting the elephant and mouse standing on opposite ends of a seesaw, where the heavier elephant naturally tilts the seesaw downward. This highlights our model’s ability to reason about relative weight, spatial arrangement, and physical dynamics implied by the instruction. In contrast, InstructPix2Pix fails to capture the core intent, producing an irrelevant result. Other methods may include the correct objects mentioned in the instruction–such as the elephant, mouse, and seesaw–but fall short in modeling their physical interactions, resulting in unrealistic or semantically inconsistent compositions. Notably, ChatGPT-4o produces an implausible outcome in which the mouse outweighs the elephant, contradicting basic physical intuition. We attribute this failure to hallucinations introduced by the language model [16].
@@ 4.4 Ablation and Analysis
Effectiveness of ReasonBrain’s Components. To evaluate the contribution of each component in ReasonBrain, we conduct ablation experiments by progressively integrating key modules. As shown in Tab. 4 and Fig. 7, the fine-grained visual features help the model capture subtle visual cues essential for reasoning. The ID Controller significantly improves performance by preserving object identity during cross-modal alignment, which is critical for accurate instruction reasoning. Additionally, the CME module enhances overall generation quality by reinforcing modality-specific semantics and providing more detailed guidance.
@@ 5 Conclusion
A Dataset Generation
The construction of our data consists of two parts. The first part (over 90% of the entire dataset) is generated following the pipeline illustrated in Fig. A1, which adopts an inverse generation strategy–deriving the source image from the target. Specifically, we first adopt the same procedure as PhyBench [29] to generate target images along with their initial instructions. Each initial instruction is then rewritten into a hypothetical form using prompt-based rewriting with GPT [1]. In parallel, we use SpaCy222https://spacy.io/ to perform named entity recognition (NER) on the initial instruction to extract candidate objects for source image generation. These candidates are passed to a diffusion model equipped with an IP-Adapter to synthesize multiple image variants. Each candidate image is subsequently evaluated by GPT, and the top-N images are selected based on a combination of GPT scores and perceptual quality metrics [11, 13, 49]. Each selected image, along with the corresponding hypothetical instruction and target image, constitutes a sample in our dataset. The second part (less than 10%) consists of story-type samples derived from EditWorld [46]. We apply the same instruction rewriting, scoring, and filtering process to refine and select high-quality samples for inclusion.
B Training Objectives of ReasonBrain
The training of ReasonBrain comprises two components: fine-tuning the MLLM and optimizing the diffusion model. Specifically, for fine-tuning the MLLM, we freeze most of its parameters and apply Low-Rank Adaptation (LoRA) [15] for efficient adaptation. The objective is defined as:
@@ B Training Objectives of ReasonBrain
where is the sampled noise and is the noisy latent at timestep . denotes the denoising network trained to predict the added noise based on the timestep and the provided visual reasoning guidance. The overall training objective is defined as the sum of the MLLM and diffusion losses: .
C More Qualitative Results

### v2

@@ 4.1 Quantitative Results
| InstructPix2Pix (Brooks et al., 2023) | 0.083 | 0.825 | 0.211 | 0.207 | 0.846 | 0.678 | 0.153 | 0.785 | 0.191 | 0.196 | 0.628 | 0.225 | 0.160 | 0.771 | 0.326 |
| MagicBrush (Zhang et al., 2023) | 0.102 | 0.844 | 0.335 | 0.228 | 0.877 | 0.725 | 0.162 | 0.802 | 0.344 | 0.209 | 0.635 | 0.359 | 0.175 | 0.790 | 0.441 |
| MGIE (Fu et al., 2024) | 0.098 | 0.802 | 0.288 | 0.213 | 0.832 | 0.685 | 0.155 | 0.761 | 0.328 | 0.201 | 0.622 | 0.322 | 0.167 | 0.754 | 0.406 |
| SmartEdit (Huang et al., 2024) | 0.118 | 0.849 | 0.602 | 0.226 | 0.881 | 0.779 | 0.165 | 0.823 | 0.385 | 0.211 | 0.655 | 0.361 | 0.180 | 0.802 | 0.532 |
| UltraEdit (Zhao et al., 2024) | 0.156 | 0.861 | 0.584 | 0.231 | 0.922 | 0.826 | 0.193 | 0.869 | 0.482 | 0.209 | 0.669 | 0.458 | 0.197 | 0.830 | 0.588 |
| PixWizard (Lin et al., 2024) | 0.161 | 0.881 | 0.481 | 0.234 | 0.951 | 0.833 | 0.132 | 0.863 | 0.393 | 0.172 | 0.697 | 0.389 | 0.175 | 0.848 | 0.524 |
@@
| InstructPix2Pix (Brooks et al., 2023) | 0.078 | 0.834 | 0.219 | 0.121 | 0.762 | 0.115 | 0.837 | 0.245 | 0.093 | 0.767 |
| MagicBrush (Zhang et al., 2023) | 0.090 | 0.838 | 0.222 | 0.100 | 0.776 | 0.123 | 0.883 | 0.261 | 0.058 | 0.871 |
| MGIE (Fu et al., 2024) | 0.083 | 0.746 | 0.231 | 0.163 | 0.594 | 0.116 | 0.745 | 0.251 | 0.162 | 0.577 |
| SmartEdit (Huang et al., 2024) | 0.092 | 0.858 | 0.274 | 0.119 | 0.771 | 0.119 | 0.895 | 0.262 | 0.094 | 0.820 |
| Emu Edit (Sheynin et al., 2024) | 0.109 | 0.859 | 0.231 | 0.094 | 0.819 | 0.135 | 0.897 | 0.261 | 0.052 | 0.879 |
| UltraEdit (Zhao et al., 2024) | 0.107 | 0.844 | 0.283 | 0.071 | 0.793 | - | 0.868 | - | 0.088 | 0.792 |
| PixWizard (Lin et al., 2024) | 0.104 | 0.845 | 0.248 | 0.069 | 0.798 | 0.124 | 0.884 | 0.265 | 0.063 | 0.876 |
@@ 4.2 Qualitative Results
As shown in Fig. 6, we visualize editing results of ReasonBrain and selected SOTA methods across four different reasoning scenarios. Our method demonstrates superior capability in executing hypothetical editing instructions by accurately reasoning about user intent, affected objects, and their plausible state transitions, while also maintaining stability in non-edited regions. For instance, in the first row (physical reasoning), our method successfully generates a physically plausible and contextually coherent scene–depicting the elephant and mouse standing on opposite ends of a seesaw, where the heavier elephant naturally tilts the seesaw downward. This highlights our model’s ability to reason about relative weight, spatial arrangement, and physical dynamics implied by the instruction. In contrast, InstructPix2Pix fails to capture the core intent, producing an irrelevant result. Other methods may include the correct objects mentioned in the instruction–such as the elephant, mouse, and seesaw–but fall short in modeling their physical interactions, resulting in unrealistic or semantically inconsistent compositions. Notably, ChatGPT-4o produces an implausible outcome in which the mouse outweighs the elephant, contradicting basic physical intuition. We attribute this failure to hallucinations introduced by the language model (Huang et al., 2025).
@@ 4.3 Ablation and Analysis
Effectiveness of ReasonBrain’s Components. To evaluate the contribution of each component in ReasonBrain, we conduct ablation experiments by progressively integrating key modules. As shown in Tab. 4 and Fig. 7, the fine-grained visual features help the model capture subtle visual cues essential for reasoning. The ID Controller significantly improves performance by preserving object identity during cross-modal alignment, which is critical for accurate instruction reasoning. Additionally, the CME module enhances overall generation quality by reinforcing modality-specific semantics and providing more detailed guidance. Furthermore, we also observe that removing any individual component leads to a performance drop across all three metrics, indicating that each module contributes meaningfully to the overall effectiveness of the proposed framework.
@@ 5 Conclusion
A Reproducibility Statement
We have already elaborated on all the models or algorithms proposed, experimental configurations, and benchmarks used in the experiments in the main body or appendix of this paper. Furthermore, we declare that the entire code used in this work will be released after acceptance.
B The Use of Large Language Models
We use large language models solely for polishing our writing, and we have conducted a careful check, taking full responsibility for all content in this work.
C Dataset Generation
The construction of our data consists of two parts. The first part (over 90% of the entire dataset) is generated following the pipeline illustrated in Fig. A1, which adopts an inverse generation strategy–deriving the source image from the target. Specifically, we first adopt the same procedure as PhyBench (Meng et al., 2024) to generate target images along with their initial instructions. Each initial instruction is then rewritten into a hypothetical form using prompt-based rewriting with GPT (Achiam et al., 2023). In parallel, we use SpaCy222https://spacy.io/ to perform named entity recognition (NER) on the initial instruction to extract candidate objects for source image generation. These candidates are passed to a diffusion model equipped with an IP-Adapter to synthesize multiple image variants. Each candidate image is subsequently evaluated by GPT, and the top-N images are selected based on a combination of GPT scores and perceptual quality metrics (He et al., 2023; Hore & Ziou, 2010; Zhang et al., 2018). Each selected image, along with the corresponding hypothetical instruction and target image, constitutes a sample in our dataset. The second part (less than 10%) consists of story-type samples derived from EditWorld (Yang et al., 2024). We apply the same instruction rewriting, scoring, and filtering process to refine and select high-quality samples for inclusion.
D Training Objectives of ReasonBrain
The training of ReasonBrain comprises two components: fine-tuning the MLLM and optimizing the diffusion model. Specifically, for fine-tuning the MLLM, we freeze most of its parameters and apply Low-Rank Adaptation (LoRA) (Hu et al., 2022) for efficient adaptation. The objective is defined as:
@@ D Training Objectives of ReasonBrain
where is the sampled noise and is the noisy latent at timestep . denotes the denoising network trained to predict the added noise based on the timestep and the provided visual reasoning guidance. The overall training objective is defined as the sum of the MLLM and diffusion losses: . Moreover, to mitigate hallucination risks, we adopted a context- and instruction-aware token selection strategy inspired by SID (Huo et al., 2025), combined with the dynamic token propagation mechanism from TAME (Tang et al., 2025) in the training process of MLLM.
E Setups: Datasets, Metrics, and Details
Datasets: We use Reason50K for both training and evaluation. Specifically, for each reasoning category, samples are randomly selected for validation, while the remaining samples are used for training. In addition, we assess the reasoning capability of ReasonBrain on two external benchmarks: ReasonEdit (Huang et al., 2024) and EditWorld (Yang et al., 2024). To evaluate generalization on conventional understanding scenarios, we further test on the MagicBrush Test Set (Zhang et al., 2023) and the Emu Edit Test Set (Sheynin et al., 2024).
Metrics: To evaluate performance under reasoning scenarios, we adopt three metrics: CLIP Score (Radford et al., 2021), MLLM Score (Yang et al., 2024), and Instruction Alignment (Ins-Align) (Huang et al., 2024). Here, CLIP Score measures the semantic similarity between the edited image and the expected output text using CLIP’s image-text embeddings. MLLM Score employs an MLLM to assess instruction-following performance. Following Yang et al. (2024), we provide the input description, editing instruction, and output description along with the edited image to Video-LLaVA. The prompt is defined as: “The input description: [object Object], the editing instruction: [object Object], and the output description: [object Object]. Please evaluate if the given edited image has been successfully edited. If yes, return 1; if not, return 0.” The final score is the average of model judgments across all samples. In addition, Ins-Align Score evaluates how well the edited image aligns with the given instruction. Following Huang et al. (2024), ten human annotators independently rated the outputs on the Reason50K dataset, and we report the average alignment score. For understanding scenarios, we adopt L1 distance, CLIP image similarity, DINO similarity, CLIP text-image similarity, and CLIP text-image direction similarity as evaluation metrics.
Implementation Details: During training, we adopt the pre-trained LLaVAv1.1-7B (Liu et al., 2024b) and QFormer (Li et al., 2023) and employ DeepSpeed (Aminabadi et al., 2022) Zero-2 to perform LoRA (Hu et al., 2021) fine-tuning, with rank and alpha of 8 and 16, respectively. Following (Huang et al., 2024; Fu et al., 2024), we expand the original LLM vocabulary with 32 new tokens, and the QFormer is composed of 6 transformer layers and 77 learnable query tokens. For the base editing model, we implement it with Flux (Labs, 2024) using FLUX.1-dev, which consists of 12B parameters. Models for other qualitative results are implemented using SD series (CompVis, 2022; SimianLuo, 2024; Rombach et al., 2022a; AI, 2022) with their original codebase. Our model is trained with a batch size of 16, and we use AdamW (Loshchilov & Hutter, 2017) optimizer to train the model with the weight decay as 1e-2 and the learning rate as 1e-3. All the experiments are conducted on 16 H20 GPUs. For a fair comparison, all baseline models are fine-tuned on the same training set used by ReasonBrain.
F More Quantitative Results
Computational costs & lightweight alternative: Tab. A1 shows the inference time of our ReasonBrain alongside all baselines. We find that our model demonstrates strong reasoning capabilities while maintaining an inference time comparable to that of existing methods. To further accelerate the model, we explored a lightweight variant, ReasonBrain-3B, by adopting a smaller pretrained MLLM. The performance comparison across all four reasoning categories between ReasonBrain-3B and our full model, ReasonBrain-7B, is illustrated in Tab. A2. We observe that, compared to ReasonBrain-7B, ReasonBrain-3B is significantly faster but exhibits only a slight performance drop. Nevertheless, it still outperforms all baselines reported in Tab. 2. This indicates that even in instruction-based reasoning scenarios, a lightweight MLLM can not only accelerate inference but also effectively identify the editing target and execute precise edits, supported by its strong reasoning ability and rich world knowledge.
| Method | Inference Time (s) |
| InstructPix2Pix | 26 |
| MagicBrush | 28 |
| MGIE | 37 |
| SmartEdit | 33 |
| UltraEdit | 28 |
| PixWizard | 35 |
| Ours | 32 |
| Model | CLIP | MLLM | Ins-Align | Inference Time (s) |
| ReasonBrain-3B | 0.238 | 0.852 | 0.822 | 24 |
| ReasonBrain-7B | 0.259 | 0.877 | 0.847 | 32 |
Multi-step editing: We extend the original one-step experiments to include 2-step and 3-step editing scenarios (multi-hop reasoning chains, e.g., ’ice melting water evaporating damp ground forming’). The results (shown in Tab. A3) indicate that the model maintains robust performance across multiple editing steps, demonstrating the effectiveness of ReasonBrain in handling sequential multi-step reasoning-based editing. We attribute this to the ID Controller module, which operates independently at each editing step, enabling effective processing of chained instructions without causing significant feature drift.
| Steps | CLIP | MLLM | Ins-Align |
| 1-step | 0.259 | 0.877 | 0.847 |
| 2-steps | 0.257 | 0.878 | 0.848 |
| 3-steps | 0.260 | 0.877 | 0.847 |
Human evaluation: We randomly select 50 images corresponding to four distinct reasoning scenarios. For each image, we obtain the results of InstructPix2Pix, MagicBrush, MGIE, SmartEdit, UltraEdit and PixWizard. Then, randomly shuffle the order of these method results. For each set of images, we ask 30 participants to independently select the three best pictures. The first one is the best picture corresponding to the text prompt (i.e., Instruct Alignment), and the second one is the picture with the highest visual quality (i.e., Image Quality). The result is shown in Tab. A4. We found that 75.71% of participants believed ReasonBrain better reflected the correct reasoning behind the instructions, and 63.55% preferred the results generated by ReasonBrain. In contrast, all baseline methods received less than 10% on both evaluations.
| Method | Instruct-Alignment (%) | Image Quality (%) |
| InstructPix2Pix | 2.62 | 2.14 |
| MagicBrush | 3.11 | 5.65 |
| MGIE | 3.10 | 6.68 |
| SmartEdit | 3.55 | 6.85 |
| UltraEdit | 6.09 | 6.11 |
| PixWizard | 5.82 | 9.02 |
| Ours | 75.71 | 63.55 |
The influences of the data generation process on the final performance: Let us briefly recall the key steps in our data generation process (see App. C and Fig. A1 for details):
Utilizing PhyBench to acquire a target image and its initial prompt;
Extracting candidate objects from the prompt (e.g., object identity, action, attributes);
Using these candidate objects to guide a T2I model to generate candidate source images;
Selecting the best candidate based on a combination of evaluation metrics;
Rewriting the initial prompt into a hypothetical instruction format.
Finally, each data sample is composed of a selected candidate image, the rewritten instruction, and the original target image. We argue that the potential sources of influence on the final performance include:
Noise in Source Image Generation (Step a): Some target images may contain distortions or semantically irrelevant content.
Missing Object Elements (Step b): Key candidate objects may be omitted during prompt decomposition.
Lack of Constraints (Step c): The T2I generation process does not explicitly enforce identity or background preservation when generating candidate source images.
Limited Evaluation Dimensions (Step d): The scoring and selection criteria may overlook fine-grained visual details such as object consistency or scene coherence.
To this end, we conduct a preliminary ablation study to assess how these factors affect model performance. The results are summarized in Tab. A5. We find that when the initial step produces low-quality target images (error image), the model’s performance drops significantly. Similarly, in the source image generation step, missing object elements (missing ID text) or ignoring ID correlations (w/o ID adapter) negatively impact performance. Finally, during the selection phase, using only a single metric (i.e., single selection) results in a considerable performance gap, highlighting the importance of applying a comprehensive set of evaluation metrics when constructing the final dataset.
| Setting | CLIP | MLLM | Ins-Align |
| Error Image | 0.104 | 0.425 | 0.226 |
| Missing ID Text | 0.125 | 0.465 | 0.233 |
| w/o ID Adapter | 0.087 | 0.312 | 0.158 |
| Single Selection | 0.172 | 0.705 | 0.505 |
| Ours | 0.259 | 0.877 | 0.847 |
Identity preservation: We followed the identity preservation metric proposed in Yang et al. (2025) and conduct an evaluation across multiple baseline methods. The results are summarized in Tab. A6. As shown, our ReasonBrain achieves the highest score, indicating that it more effectively retains core elements from the source image while following the instruction. This demonstrates that our model trained on Reason50K is actually performing reasoning and not just learning a shortcut.
| Method | Identity Preservation |
| InstructPix2Pix | 5.13 |
| MagicBrush | 7.52 |
| MGIE | 7.63 |
| SmartEdit | 8.02 |
| UltraEdit | 5.68 |
| PixWizard | 8.55 |
| Ours | 9.72 |
Effectiveness of CME module: To validate the effectiveness of the bidirectional information interaction in our proposed CME module, we conducted two comparative experiments. In Exp 1, we remove the CME module entirely and directly feed the feature output from QFormer into the diffusion model. This ablation study is designed to evaluate the effectiveness of the information interaction introduced by the CME module. In Exp 2, we aim to assess the necessity of bidirectional information interaction. Specifically, we retain only the cross-attention block on the image feature branch, discarding all other components of the CME module. As a result, the textual features from QFormer are applied to the image features in a unidirectional manner. The results are presented in Tab. A7, demonstrating the marginal gains brought by each component and highlighting the overall benefit of the complete CME design.
| Exp ID | Plain | Simple CA | CME | CLIP | MLLM | Ins-Align |
| 1 | ✓ | 0.239 | 0.833 | 0.758 | ||
| 2 | ✓ | 0.241 | 0.841 | 0.766 | ||
| Ours | ✓ | 0.259 | 0.877 | 0.847 |
Testing on different mapping settings: We conduct an additional ablation study by extending our original one-to-one mapping to one-to-two and one-to-three mappings. The results are shown in Tab. A8, we find that incorporating multiple targets (i.e., one-to-two/one-to-three mappings) led to significant performance degradation across all metrics. These settings introduced training instability due to conflicting optimization signals from multiple targets for the same input. In particular, Ins-Align and Identity Preservation dropped significantly, indicating that using multiple targets for the same input makes it difficult to retain key elements from the source image–further emphasizing the importance of maintaining a one-to-one structure in our task.
| Setting | CLIP | MLLM | Ins-Align | Identity Preservation |
| Ours (one-to-one) | 0.259 | 0.877 | 0.847 | 9.72 |
| one-to-two | 0.152 | 0.680 | 0.315 | 4.25 |
| one-to-three | 0.088 | 0.258 | 0.204 | 1.66 |
G Limitations
Limitations on Real-world Data Collection. The construction of our dataset Reason50K follows the common practice in the image editing community (e.g., EmuEdit (Sheynin et al., 2024) and GPT-Image-Edit (Wang et al., 2025)), relying primarily on synthetic data. While synthetic datasets provide semantic clarity and high visual quality, they may not fully capture the complexities, artifacts, and temporal dynamics inherent in real-world video data. Collecting high-quality datasets from real-world video sources remains particularly challenging due to the scarcity of suitable image pairs with explicit reasoning logic and the absence of standardized benchmarks or evaluation protocols for reasoning-based edits. Moreover, prior efforts such as EditWorld (Yang et al., 2024) show that frames extracted from videos often suffer from low resolution and poor aesthetics, making them suboptimal for training high-fidelity models. Thus, addressing this gap by curating reasoning-aligned, high-quality video benchmarks constitutes an important direction for our future work.
H More Qualitative Results